# Relatives of the Hermitian curve

### v2

@@ Abstract
with $A\in GL(3,\mathbb{F}_{q})$, and study their basic properties together with various examples. One of the basic properties is that the number of $\mathbb{F}_{q}$-points of any relative of the Hermitian curve of degree $\sqrt{q}+1$ is congruent to $1$ modulo $\sqrt{q}$.
@@
1 Our object
Throughout this paper, $q$ denotes an even power of a prime number, so $\sqrt{q}$ is also a power of the prime number.
A Hermitian curve is a plane curve defined by
$(x^{\sqrt{q}},y^{\sqrt{q}},z^{\sqrt{q}})A\begin{pmatrix}x\\ y\\ z\end{pmatrix}=0,$ (1)
where $A\in GL(3,\mathbb{F}_{q})$ satisfying the Hermitian condition ${}^{t}\!A^{(\sqrt{q})}=A$. Here the symbol $A^{(\sqrt{q})}$ means taking $\sqrt{q}$-th power on each entry of $A$.
@@ A Hermitian curve is a plane curve defined by
All Hermitian-relative curves, including Hermitian curves themselves, are projectively equivalent each other over the algebraic closure $\overline{\mathbb{F}}_{q}$ of $\mathbb{F}_{q}$, which is due to Pardini [7, Proof of Proposition 3.7]. We will give a proof of this fact in the next section for the convenience of readers.
@@
The proof of this fact is also given in the following section.
Since the equation (1) is homogeneous, the curve $C_{A}$ depends only on the image $\overline{A}$ of $A$ under the canonical map $GL(3,\mathbb{F}_{q})\to PGL(3,\mathbb{F}_{q})$. By abuse of notation, $\overline{A}$ will be frequently denoted by $A$, namely an element of $GL(3,\mathbb{F}_{q})$ and the image of the element in $PGL(3,\mathbb{F}_{q})$ will be used interchangeably. So the equation $A=B$ in $PGL(3,\mathbb{F}_{q})$ means that there exists an element $\lambda\in\mathbb{F}_{q}^{\ast}$ so that $A=\lambda B$ in $GL(3,\mathbb{F}_{q})$.
Remark 1.1
If $\overline{A}\in PGL(3,\mathbb{F}_{q})$ satisfies the Hermitian condition ${}^{t}\overline{A}^{(\sqrt{q})}=\overline{A}$, then we can choose an element $A\in GL(3,\mathbb{F}_{q})$ satisfying the Hermitian condition whose image in $PGL(3,\mathbb{F}_{q})$ is the assigned $\overline{A}$.
@@
Let $A$ be any source of $\overline{A}$. Then there is an element $\lambda\in\mathbb{F}_{q}^{\ast}$ such that ${}^{t}A^{(\sqrt{q})}=\lambda A$. Taking $\sqrt{q}$-th power and transpose of the both side, $A=\lambda^{\sqrt{q}}\left({}^{t}\!A^{(\sqrt{q})}\right)$. So $A=\lambda^{\sqrt{q}+1}A$, which implies $\lambda^{\sqrt{q}+1}=1$. Therefore there is an element $\rho\in\mathbb{F}_{q}$ so that $\rho^{1-\sqrt{q}}=\lambda$ (which is a special case of Hilbert’s Theorem 90). Then $\overline{\rho A}=\overline{A}$ and $\rho A$ satisfies the Hermitian condition. ∎
Annotation
Although a homogeneous coordinates of a point has an ambiguity by nonzero multiple, hereafter, for any $\mathbb{F}_{q}$-point, we understand all values of the coordinates are elements of $\mathbb{F}_{q}$.
2 Basic properties
@@
$\begin{pmatrix}x\\ y\\ z\end{pmatrix}\mapsto T\begin{pmatrix}x\\ y\\ z\end{pmatrix}$
@@
$(x^{\sqrt{q}},y^{\sqrt{q}},z^{\sqrt{q}}){}^{t}T^{(\sqrt{q})}AT\begin{pmatrix}x\\ y\\ z\end{pmatrix}.$
The properties mentioned above hold true after any field extension $\mathbb{F}_{q}$.
For given two matrices $A,B\in GL(3,\mathbb{F}_{q})$ (or in $PGL(3,\mathbb{F}_{q})$), we say that $A$ and $B$ are equivalent if two curves $C_{A}$ and $C_{B}$ are projectively equivalent over $\mathbb{F}_{q}$. This equivalence relation is denoted by $A\sim B$. In other words, $A\sim B$ if there is a matrix $T\in GL(3,\mathbb{F}_{q})$ and $\lambda\in\mathbb{F}_{q}^{\ast}$ such that $\lambda B={}^{t}T^{(\sqrt{q})}AT$ (or $B={}^{t}T^{(\sqrt{q})}AT$ in $PGL(3,\mathbb{F}_{q})$).
Remark 2.1
@@
Proposition 2.2
For $A\in GL(3,\mathbb{F}_{q})$, the dual curve of $C_{A}$ is projectively equivalent to $C_{A^{(\sqrt{q})}}$ over $\mathbb{F}_{q}$.
Proof.
$(u,v,w)=(x^{\sqrt{q}},y^{\sqrt{q}},z^{\sqrt{q}})A$ (3)
which can be written
by (3). On the other hand, since
(${}^{t}\!\left({}^{t}\!A\right)^{(\sqrt{q})}\left({}^{t}\!A^{-1}\right){}^{t}\!A=A^{(\sqrt{q})},$
${}^{t}\!A^{-1}\sim A^{(\sqrt{q})}$. This completes the proof. ∎
Next is the first main purpose of this section.
Theorem 2.3
@@
$f(x,y,z)=(x^{\sqrt{q}},y^{\sqrt{q}},z^{\sqrt{q}})A\begin{pmatrix}x\\ y\\ z\end{pmatrix}.$
@@
$=(x^{\sqrt{q}},y^{\sqrt{q}},z^{\sqrt{q}}){}^{t}T^{(\sqrt{q})}T\begin{pmatrix}x\\ y\\ z\end{pmatrix}.$
Note that each entry of ${}^{t}T^{(\sqrt{q})}T$ gives a coefficient of distinct monomial of $f(x,y,z)$. So ${}^{t}T^{(\sqrt{q})}T\in PGL(3,\mathbb{F}_{q})$.
@@
$\Phi_{A}:PGL(3,\overline{\mathbb{F}}_{q})\ni T\mapsto{}^{t}T^{(\sqrt{q})}AT\in PGL(3,\overline{\mathbb{F}}_{q}).$
@@
The automorphism group of $C_{A}$ over $\overline{\mathbb{F}}_{q}$ is finite if $\sqrt{q}>2$, because the genus of $C_{A}$ is grater than $1$. When $\sqrt{q}=2$, $C_{A}$ is an elliptic curve over $\overline{\mathbb{F}}_{q}$ in $\mathbb{P}^{2}$. The linear automorphism group of any elliptic curve in $\mathbb{P}^{2}$ is also finite. Therefore the image of $\Phi_{A}$ is a constructible set (which is a theorem of Chevalley [6, Corollary 2 in I §8]) of dimension $8$ in the target space $PGL(3,\overline{\mathbb{F}}_{q})$. Therefore two images of $\Phi_{A}$ and $\Phi_{B}$ meet in $PGL(3,\overline{\mathbb{F}}_{q})$ for any two $A,B\in PGL(3,\mathbb{F}_{q})$, that is, there is an element $T\in PGL(3,\overline{\mathbb{F}}_{q})$ such that ${}^{t}T^{(\sqrt{q})}AT=B$. In particular, if one takes as $B=I_{3}$, one gets a projective transformation over $\overline{\mathbb{F}}_{q}$ which sends $C_{A}$ to a Hermitian curve. ∎
@@
This characterization comes from [4, Theorem 6.1, and Theorem 3.5] together with Theorem 2.3
The following is the second main purpose of this section.
Theorem 2.5
@@
In particular, the set $C(\mathbb{F}_{q})$ of $\mathbb{F}_{q}$-points on $C$ is not empty.
@@
(As for this classical theory, one may consult [9, Chapter 5].)
@@
Since $\sqrt{q}$ is an integer by our assumption, $\sum_{j=1}^{2g}\eta_{j}$ is a rational number. Therefore it is a rational integer. Hence
@@
The additional statement is somewhat classical [3, 8, 5]. For the first statement, when $q=4$, there is nothing new from Theorem 2.5. Hence we assume that $q>4$. Let $C=C_{A}$, and $(x_{0},y_{0},z_{0})\in C(\mathbb{F}_{q})$. Then $(x_{0}^{\sqrt{q}},y_{0}^{\sqrt{q}},z_{0}^{\sqrt{q}})A\begin{pmatrix}x_{0}\\ y_{0}\\ z_{0}\end{pmatrix}=0.$ Taking $\sqrt{q}$-th power, and noting $x_{0}^{q}=x_{0},y_{0}^{q}=y_{0},z_{0}^{q}=z_{0}$, we know that $(x_{0},y_{0},z_{0})\in C_{{}^{t}\!A^{(\sqrt{q})}}$. If $C$ is not Hermitian, $C_{A}$ and $C_{{}^{t}\!A^{(\sqrt{q})}}$ are distinct irreducible curves. Hence the intersection of these two curves consists of at most $(\sqrt{q}+1)^{2}$ points, which gives an upper bound for $N_{q}(C)$. ∎
3 Example I - diagonal case
Remaining three sections, selected examples are explained; however, the division of sections is rather superficial (cf. Example 3.4 and Example 4.1). In order to compute examples, some notes will be explained. The trace $\mathbb{F}_{q}\ni x\mapsto x^{\sqrt{q}}+x\in\mathbb{F}_{\sqrt{q}}$ and the norm $\mathbb{F}_{q}^{\ast}\ni x\mapsto x^{\sqrt{q}+1}\in\mathbb{F}_{\sqrt{q}}^{\ast}$ are indicated by $\operatorname{Tr}$ and $\operatorname{Nm}$ respectively.
Lemma 3.1

### v3

@@ Abstract
with $A\in GL(3,\mathbb{F}_{q})$, and study their basic properties, one of which is that the number of $\mathbb{F}_{q}$-points of any relative of the Hermitian curve of degree $\sqrt{q}+1$ is congruent to $1$ modulo $\sqrt{q}$.
In the latter part of this paper, we classify those curves having two or more rational inflexions.
@@
1 Introduction
Throughout this paper, $q$ denotes an even power of a prime number $p$. So $\sqrt{q}$ is also a power of $p$. For $A\in GL(3,\mathbb{F}_{q})$, $A^{\ast}$ denotes the matrix ${}^{t}\!A^{(\sqrt{q})}$, where the superscript $(\sqrt{q})$ indicates taking entry-wise $\sqrt{q}$-th power, and the superscripted prefix $t$ taking transpose.
The Hermitian curve is a plane curve defined by
$(x^{\sqrt{q}},y^{\sqrt{q}},z^{\sqrt{q}})A\,{}^{t}\!(x,y,z)=0,$ (1)
with the condition $A=A^{\ast}$
In this case, we say that $A$ satisfies the Hermitian condition.
@@ 1 Introduction
All Hermitian-relative curves, including Hermitian curves themselves, are projectively equivalent each other over the algebraic closure $\overline{\mathbb{F}}_{q}$ of $\mathbb{F}_{q}$, which is due to Pardini [5, Proof of Proposition 3.7]. We will give a proof of this fact in the next section for the convenience of readers.
@@
The proof of this fact is also given in the following section. Since the equation (1) is homogeneous, the curve $C_{A}$ depends only on the image $\overline{A}$ of $A$ under the canonical map $GL(3,\mathbb{F}_{q})\to PGL(3,\mathbb{F}_{q})$. By abuse of notation, $\overline{A}$ will be frequently denoted by $A$, namely an element of $GL(3,\mathbb{F}_{q})$ and its image in $PGL(3,\mathbb{F}_{q})$ will be used interchangeably. So the equation $A=B$ in $PGL(3,\mathbb{F}_{q})$ means that there exists an element $\lambda\in\mathbb{F}_{q}^{\ast}$ so that $A=\lambda B$ in $GL(3,\mathbb{F}_{q})$.
Remark 1.1.
If $\overline{A}\in PGL(3,\mathbb{F}_{q})$ satisfies the Hermitian condition, that is to say, $\overline{A^{\ast}}=\overline{A}$, then we can choose an element $A\in GL(3,\mathbb{F}_{q})$ satisfying the Hermitian condition whose image in $PGL(3,\mathbb{F}_{q})$ is the assigned $\overline{A}$.
@@
Let $A$ be any source of $\overline{A}$. Then there is an element $\lambda\in\mathbb{F}_{q}^{\ast}$ such that $A^{\ast}=\lambda A$. Taking $\ast$-operation on the both side, $A=\lambda^{\sqrt{q}}A^{\ast}$. So $A=\lambda^{\sqrt{q}+1}A$, which implies $\lambda^{\sqrt{q}+1}=1$. Therefore there is an element $\rho\in\mathbb{F}_{q}$ so that $\rho^{1-\sqrt{q}}=\lambda$ (which is a special case of Hilbert’s Theorem 90). Then $\overline{\rho A}=\overline{A}$ and $\rho A$ satisfies the Hermitian condition. ∎
@@
$\begin{pmatrix}x\\ y\\ z\end{pmatrix}\mapsto T^{-1}\begin{pmatrix}x\\ y\\ z\end{pmatrix}$
@@
$(x^{\sqrt{q}},y^{\sqrt{q}},z^{\sqrt{q}})T^{\ast}AT\begin{pmatrix}x\\ y\\ z\end{pmatrix}.$
In other words, two relatives of the Hermitian curve $C_{A}$ and $C_{B}$, are projectively equivalent over $\mathbb{F}_{q}$ if and only if there is a certain matrix $T\in PGL(3,\mathbb{F}_{q})$ such that $B=T^{\ast}AT$ in $PGL(3,\mathbb{F}_{q})$. This is sometimes denoted by $A\sim B$.
Remark 1.2.
@@
The properties mentioned above hold true after any field extension $\mathbb{F}_{q}$. As is explained in the next section, the tangent line at a point of $C_{A}$ contacts of multiplicity $\sqrt{q}$ in general, but at some special points of multiplicity $\sqrt{q}+1$. In the letter case, the point is called an inflexion. In Section 3, we give the classification up to $PGL(3,\mathbb{F}_{q})$ of the curves $C_{A}$ with two or more inflexions defined over $\mathbb{F}_{q}$. In the classification, three types appear and it can be summarized in Table 1.
Table 1: Classification
$N_{q}(C_{A})$ inflexions classes
(a) $\sqrt{q}^{3}+1$ $\sqrt{q}^{3}+1$ 1
(b) $\sqrt{q}+1$ $\sqrt{q}+1$ $\sqrt{q}$
(c) $q+1$ $2$ $\frac{1}{2}(\sqrt{q}+1)(\sqrt{q}-2)$
In the table, the first column indicates the label of each type which agrees with the item in Theorem 3.1, the second the number of $\mathbb{F}_{q}$-points of $C_{A}$ belongs the assigned type, the third the number of inflexions, and the fourth is the number of $\mathbb{F}_{q}$-equivalent classes in each type.
Annotation.
$\bullet$ Although a homogeneous coordinates of a point has an ambiguity by nonzero multiple, we understand all values of the coordinates of any $\mathbb{F}_{q}$-point are elements in $\mathbb{F}_{q}$. Moreover, we use two words “$\mathbb{F}_{q}$-point” and “rational point” interchangeably.
$\bullet$ The set of $\mathbb{F}_{q}$-points of a curve $C$ is denoted by $C(\mathbb{F}_{q})$.
$\bullet$ For a field, the multiplicative group of nonzero elements is marked by the superscript $\ast$.
$\bullet$ The trace map from $\mathbb{F}_{q}$ to $\mathbb{F}_{\sqrt{q}}$, that is, $\mathbb{F}_{q}\ni\alpha\mapsto\alpha+\alpha^{\sqrt{q}}\in\mathbb{F}_{\sqrt{q}},$ and the norm map $\mathbb{F}_{q}^{\ast}$ to $\mathbb{F}_{\sqrt{q}}^{\ast}$, that is, $\mathbb{F}_{q}^{\ast}\ni\alpha\mapsto\alpha^{\sqrt{q}+1}\in\mathbb{F}_{\sqrt{q}}^{\ast}$ are simply denoted by $\operatorname{Tr}$ and $\operatorname{Nm}$ respectively.
$\bullet$ Unless we specify the field of definition, ‘projectively equivalent’ means ‘projectively equivalent over $\mathbb{F}_{q}$’.
$\bullet$ Sometimes a curve defined by $g=0$ will be indicated by $\{g=0\}$.
2 Basic properties
We shall start this section with Pardini’s observation [5], which is our motivation to study the Hermitian-relative curve.
Theorem 2.1.
@@ 2 Basic properties
$f(x,y,z)=(x^{\sqrt{q}},y^{\sqrt{q}},z^{\sqrt{q}})A{}^{t}\!(x,y,z).$
@@
$=(x^{\sqrt{q}},y^{\sqrt{q}},z^{\sqrt{q}})T^{\ast}T\begin{pmatrix}x\\ y\\ z\end{pmatrix}.$
Note that each entry of $T^{\ast}T$ gives a coefficient of distinct monomial of $f(x,y,z)$. So $T^{\ast}T\in PGL(3,\mathbb{F}_{q})$.
@@
$\Phi_{A}:PGL(3,\overline{\mathbb{F}}_{q})\ni T\mapsto T^{\ast}AT\in PGL(3,\overline{\mathbb{F}}_{q}).$
@@
The automorphism group of $C_{A}$ over $\overline{\mathbb{F}}_{q}$ is finite if $\sqrt{q}>2$, because the genus of $C_{A}$ is grater than $1$. When $\sqrt{q}=2$, $C_{A}$ is an elliptic curve over $\overline{\mathbb{F}}_{q}$ in $\mathbb{P}^{2}$. The linear automorphism group of any elliptic curve in $\mathbb{P}^{2}$ is also finite. Therefore the image of $\Phi_{A}$ is a constructible set (which is a theorem of Chevalley [4, Corollary 2 in I §8]) of dimension $8$ in the target space $PGL(3,\overline{\mathbb{F}}_{q})$. Therefore two images of $\Phi_{A}$ and $\Phi_{B}$ meet in $PGL(3,\overline{\mathbb{F}}_{q})$ for any two $A,B\in PGL(3,\mathbb{F}_{q})$, that is, there is an element $T\in PGL(3,\overline{\mathbb{F}}_{q})$ such that $T^{\ast}AT=B$. In particular, if one takes as $B=I_{3}$, one gets a projective transformation over $\overline{\mathbb{F}}_{q}$ which sends $C_{A}$ to a Hermitian curve. ∎
@@
This characterization comes from [2, Theorem 6.1, and Theorem 3.5] together with Theorem 2.1
The following is the main purpose of this section.
Theorem 2.3.
@@
In particular, the set $C(\mathbb{F}_{q})$ of $\mathbb{F}_{q}$-points is not empty.
@@
(As for this classical theory, one may consult [7, Chapter 5].)
@@
Since $\sqrt{q}$ is a rational integer by our assumption, $\sum_{j=1}^{2g}\eta_{j}$ is a rational number. Therefore it is a rational integer. Hence
@@
The additional statement is somewhat classical [1, 6, 3]. For the first statement, when $q=4$, there is nothing new from Theorem 2.3. Hence we assume that $q>4$. Let $C=C_{A}$, and $(x_{0},y_{0},z_{0})\in C(\mathbb{F}_{q})$. Then $(x_{0}^{\sqrt{q}},y_{0}^{\sqrt{q}},z_{0}^{\sqrt{q}})A{}^{t}\!(x_{0},y_{0},z_{0})=0.$ Taking $\sqrt{q}$-th power, and noting $x_{0}^{q}=x_{0},y_{0}^{q}=y_{0},z_{0}^{q}=z_{0}$, we know that $(x_{0},y_{0},z_{0})\in C_{A^{\ast}}$. If $C$ is not Hermitian, $C_{A}$ and $C_{A^{\ast}}$ are distinct irreducible curves. Hence the intersection of these two curves consists of at most $(\sqrt{q}+1)^{2}$ points, which gives an upper bound for $N_{q}(C)$. ∎
Before closing this section, we shall digress a little to explain a couple of geometric properties of relatives of the Hermitian curve.
Proposition 2.5.
For $A\in GL(3,\mathbb{F}_{q})$, the dual curve of $C_{A}$ is $C_{{}^{t}\!A^{-1}}$.
Proof.
$(u,v,w)=(x^{\sqrt{q}},y^{\sqrt{q}},z^{\sqrt{q}})A$ (4)
which can be written
by (4). This completes the proof. ∎
For a Hermitian-relative curve $C_{A}$, the curve $C_{A^{\ast}}$ is called the mirror curve of $C_{A}$. Of course, the mirror curve of $C_{A^{\ast}}$ is the original one $C_{A}$, and $C_{A}$ is Hermitian if and only if $C_{A^{\ast}}=C_{A}$.
Proposition 2.6.
For a mirror pair $(C_{A},C_{A^{\ast}})$, the following properties hold.
  1. (1)
    $C_{A}(\mathbb{F}_{q})=C_{A^{\ast}}(\mathbb{F}_{q})$.
  2. (2)
    For $P\in C_{A}(\mathbb{F}_{q})=C_{A^{\ast}}(\mathbb{F}_{q})$, $P$ is an inflexion of $C_{A}$ if and only if an inflexion of $C_{A^{\ast}}$. In this case, $T_{P}(C_{A})=T_{P}(C_{A}^{\ast})$.
  3. (3)
    For a non-inflexion $P\in C_{A}(\mathbb{F}_{q})$, then there is a point $P^{\prime}\in C_{A}(\mathbb{F}_{q})$ so that $C_{A}.T_{P}C_{A}=\sqrt{q}P+P^{\prime}$. In this case, $C_{A^{\ast}}.T_{P^{\prime}}C_{A^{\ast}}=\sqrt{q}P^{\prime}+P$. In particular $P^{\prime}$ is a non-inflexion of $C_{A^{\ast}}$.
Proof.
If $C_{A}$ is Hermitian, then the assertions (1) and (2) are trivially true, and (3) is nonsense. So $C_{A}$ may be assumed to be non-Hermitian.
We already saw the assertion (1) in the proof of Corollary 2.4.
(2) Since $(T^{\ast}AT)^{\ast}=T^{\ast}A^{\ast}T$ for any $T\in PGL(3,\mathbb{F}_{q})$, we may assume that $P=(0,0,1)$ and $T_{P}C_{A}=\{y=0\}$. Let $A=(a_{ij})$. Since $P\in C_{A}$, $a_{33}=0$. Hence an affine equation of $C_{A}$ on $\{z\neq 0\}$ is given by
$a_{11}x^{\sqrt{q}+1}+(a_{12}y+a_{13})x^{\sqrt{q}}+(a_{21}x+a_{22}y+a_{23})y^{\sqrt{q}}+a_{31}x+a_{32}y=0.$
Since $T_{P}C_{A}=\{y=0\}$ and the origin $P$ is an inflexion of $C_{A}$, we have $a_{31}=0$, $a_{13}=0$ and $a_{11}\neq 0$. Hence we may assume that $a_{11}=1$. To sum up,
$A=\begin{pmatrix}1&a_{12}&0\\ a_{21}&a_{22}&a_{23}\\ 0&a_{32}&0\end{pmatrix}.$
Hence
$A^{\ast}=\begin{pmatrix}1&a_{21}^{\sqrt{q}}&0\\ a_{12}^{\sqrt{q}}&a_{22}^{\sqrt{q}}&a_{32}^{\sqrt{q}}\\ 0&a_{23}^{\sqrt{q}}&0\end{pmatrix},$
in other words, an affine equation of $C_{A^{\ast}}$ is
$x^{\sqrt{q}+1}+a_{21}^{\sqrt{q}}yx^{\sqrt{q}}+(a_{12}^{\sqrt{q}}x+a_{22}^{\sqrt{q}}y+a_{32}^{\sqrt{q}})y^{\sqrt{q}}+a_{32}^{\sqrt{q}}y=0.$
This shows that the origin is also an inflexion of $C_{A^{\ast}}$ with tangent line $\{y=0\}$.
(3) By Corollary 2.2, $C_{A}.T_{P}C_{A}=\sqrt{q}P+P^{\prime}$ for some point $P^{\prime}$. Since two divisors $C_{A}.T_{P}C_{A}$ and $\sqrt{q}P$ are defined over $\mathbb{F}_{q}$, so is the residual one $P^{\prime}$. Hence $P^{\prime}\in C_{A}(\mathbb{F}_{q})=C_{A^{\ast}}(\mathbb{F}_{q})$. Let $P=(x_{0},y_{0},z_{0})$ and $P^{\prime}=(x_{1},y_{1},z_{1})$. Then
$(x_{0}^{\sqrt{q}},y_{0}^{\sqrt{q}},z_{0}^{\sqrt{q}})A{}^{t}\!(x_{1},y_{1},z_{1})$
by Remark 2. Hence
$(x_{1}^{\sqrt{q}},y_{1}^{\sqrt{q}},z_{1}^{\sqrt{q}})A^{\ast}{}^{t}\!(x_{0},y_{0},z_{0}),$
which means $P\in C_{A^{\ast}}.T_{P^{\prime}}C_{A^{\ast}}$, and hence $C_{A^{\ast}}.T_{P^{\prime}}C_{A^{\ast}}=\sqrt{q}P^{\prime}+P.$
3 Hermitian-relative curves with two or more rational inflexions
The purpose of this section is to give the classification of the Hermitian-relative curves having two or more rational inflexions, which is summarized in the following theorem.
Theorem 3.1.
Let $C$ be a relative of the Hermitian curve. Suppose that $C$ has at least two rational inflexions. Then $C$ is projectively equivalent to
$x^{\sqrt{q}}y+\omega xy^{\sqrt{q}}+z^{\sqrt{q}+1}=0$ (5)
[…]
  1. (a)
[…]
  2. (b)
[…]
    $x^{\sqrt{q}+1}+y^{\sqrt{q}+1}+\eta z^{\sqrt{q}+1}=0$ (6)
[…]
    For two curves $x^{\sqrt{q}+1}+y^{\sqrt{q}+1}+\eta z^{\sqrt{q}+1}=0$ and $x^{\sqrt{q}+1}+y^{\sqrt{q}+1}+\eta^{\prime}z^{\sqrt{q}+1}=0$ with $\eta,\,\eta^{\prime}\in\mathbb{F}_{q}^{\ast}\setminus\mathbb{F}_{\sqrt{q}}^{\ast}.$ are projectively equivalent to each other if and only if $\eta\eta^{\prime-1}\in\mathbb{F}_{\sqrt{q}}^{\ast}$. Therefore, there are exactly $\sqrt{q}$ non-equivalent Hermitian-relative curves, each of which has $\sqrt{q}+1$ rational points and all of them are inflexions.
  3. (c)
[…]
3.1 Some equations
Since any trace map and norm map between finite fields are surjective, it is easy to show Hilbert’s Theorem 90 in our case:
Lemma 3.2.